OpenStudy (anonymous):

Check my answers + Will medal and fan! A restaurant did a survey among 100 customers to find their food preferences. The customers were asked about their preferences for pasta or rice. Out of the total 30 people who liked pasta, 10 liked rice also. There were 40 people who liked rice. Part A: Summarize the data by writing the values that the alphabets A to I in the table below represent. (5 points)

OpenStudy (usukidoll):

so there's 100 customers in total 40 liked rice only 30 liked pasta, but 10 of those guys liked rice too so that one is a split (20 pasta, 10 liked rice and pasta) but just adding 40+30+10 = 80?! where is the remaining 20 people? unless they don't like rice and pasta?

OpenStudy (anonymous):

There were 20 people who liked pasta (and disliked rice), and there were 40 people who liked rice. 20 + 40 = 60, so I'm assuming the remaining 40 didn't like either pasta or rice

OpenStudy (anonymous):

Because there's 30 people who liked pasta, but 10 of those people ALSO liked rice. The total number of people who liked rice is 40. So you'd take away the 10, because those people are already counted in the number 40. Then you'd have 20 + 40 = 60

OpenStudy (usukidoll):

the first column you have is correct. 30 people love pasta, but there are 10 people who love both pasta and rice.. so it splits to 10 people who like pasta and rice 20 people who like pasta, but not rice. so our total is 30

OpenStudy (anonymous):

Yeah, but since there's 40 people who like rice, the total of the row has to equal 40. Since 10 people like rice and pasta, that means that 30 people have to like rice and don't like pasta. But my total for G does equal 30.

OpenStudy (anonymous):

I appreciate all the help, but I just submitted my work :P It's 11:52pm here and I had to submit it before 12am. Thank you for your time and help, though @UsukiDoll :)

OpenStudy (usukidoll):

oh yeah xD there are 40 people who love rice... but the problem is 10 of them love rice and pasta.. so we have to take those guys out 40-10 = 30 people who love rice. actually I was going vertically I was doing A D G First because it was easier

OpenStudy (anonymous):

Yeah :P It's no problem, I've been doing these for a month and I'm still confused by them sometimes.

OpenStudy (usukidoll):

then I reached your B and C

OpenStudy (usukidoll):

10 30 40 20 E F 30 H I

OpenStudy (usukidoll):

E is the critical letter because without that we can't have H and F and without F we can't do I

OpenStudy (anonymous):

10 30 40 20 40 60 30 70 100

OpenStudy (usukidoll):

awww D": I was getting there xD. but the rows and columns do match up nicely horizontal and vertical

OpenStudy (anonymous):

Well thank god for that :D

OpenStudy (usukidoll):

this was kind of fun though xD

OpenStudy (usukidoll):

I'm still trying to figure out E which is the critical part.. once that's been done everything can be added up nicely :)

OpenStudy (usukidoll):

from what I got so far the first column you have is correct. 30 people love pasta, but there are 10 people who love both pasta and rice.. so it splits to 10 people who like pasta and rice 20 people who like pasta, but not rice. so our total is 30 B was our 30 because originally 40 people liked rice, but 10 of them like rice and pasta, so that group is thrown out 40-10 = 30 10 30 40 20 E F 30 H I

OpenStudy (usukidoll):

so all we need is the neysayers of rice and pasta.

OpenStudy (anonymous):

E kinda goes back to what we said in the beginning. 20 people liked pasta (but didn't like rice) and 40 people liked rice. 20 + 40 = 60. 40 is the remainder, so that means 40 people didn't like either pasta or rice

OpenStudy (usukidoll):

OMG XD correct 20 people liked pasta ONLY 40 peole liked rice ONLY adding those picky eaters is 60 100-60=40 !

OpenStudy (usukidoll):

10 30 40 20 40 F 30 H I

OpenStudy (anonymous):

Yup! :D

OpenStudy (usukidoll):

then we just ADD :D

OpenStudy (anonymous):

30 + 40 = 70 and 40 + 60 = 100 :D

OpenStudy (usukidoll):

10 30 40 20 40 60 30 70 100

OpenStudy (anonymous):

RIGHT! ^__^

OpenStudy (usukidoll):

that was neat... it's been a while since I last had that problem.. but that was fun like a puzzle XD

OpenStudy (anonymous):

Yeah they're definitely confusing at first but you're right, they're like a puzzle :P It is kinda fun.

OpenStudy (usukidoll):

I started with the easier to find the answer entries... which was the entire first column then just staring at the 40 love rice and 10 love rice and pasta so 40-10= 30 true rice lovers

OpenStudy (usukidoll):

rice and pasta was 10 30 true rice lovers and I ended up with 10 30 40 20 30 then there pasta lovers only from the total of 30 = 20 (pasta lovers) +10 (rice and pasta)

OpenStudy (anonymous):

Very good! ^-^

OpenStudy (usukidoll):

E was the picky eaters 20 pasta 40 rice 60 in total 100 customers - 60 = 40

OpenStudy (usukidoll):

so after having E everything else became easy

OpenStudy (anonymous):

Like a puzzle :B Once you have the right amount of pieces, everything else becomes easy :)